# Scaling limit of heavy-tailed nearly unstable cumulative INAR($\infty$) processes and rough fractional diffusions

### v6

@@ 5.2 Proof of Theorem 1
We now turn to $T_{3}$. Remark that $M_{s}=\sqrt{\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}}Z_{s/T}^{T}$, so for sufficiently large $T$,
@@
Then we can imitate the Proof of Theorem 3.1 from Jaisson and Rosenbaum (2016), and obtain the same result, $T_{3}$ converges to
$\frac{1}{\nu^{*}\lambda}\int_{0}^{t}\lambda(t-s)^{\alpha-1}E_{\alpha,\alpha}(-\lambda(t-s)^{\alpha})Z_{s}ds.$
Since $Z$ is a continuous martingale, the fact that $Z_{t}=B_{Y_{t}}$ is a consequence of the Dambis-Dubin-Schwarz theorem: see, for example, Theorem V-1.65 in Revuz and Yor (2013).
5.3 Proof of Corollary 2
The proof extends the arguments used to establish Theorem 1. The core of the argument lies in analyzing the convergence of the scaled process $Y_{t}^{T}$ under the new time-dependent baseline intensity $\hat{\mu}_{T}(n)$.
@@ 5.3 Proof of Corollary 2
Let us denote the deterministic mean as $\bar{\Lambda}_{t}^{T}:=\mathbb{E}[Y_{t}^{T}]$. Using the expression for $\mathbb{E}[N_{n}^{T}]$ from Lemma 9 and applying the scaling factor, we can decompose $\bar{\Lambda}_{t}^{T}$ into several terms. After substituting the definition of $\hat{\mu}_{T}(n)$, the expression for $\mathbb{E}[N_{n}^{T}]$ becomes:
@@
The centered stochastic part, $Y_{t}^{T}-\bar{\Lambda}_{t}^{T}$, is obtained by scaling the expression from Equation (20) in Lemma 9. We denote this scaled stochastic part as $T_{3}^{\text{stoch}}$:
@@
  • By Assumption 2, it is straightforward to show that $\lim_{T\to\infty}T_{1}=0$.
  • For $T_{2}$, we use the same Riemann sum approximation argument as in the proof of Theorem 1, where the sum $\sum_{s=1}^{n-1}sA_{n-s}^{T}$ corresponds to the integral of the CDF. In the limit, the sum of the deterministic parts converges:
@@
  • The convergence of the scaled stochastic part, $T_{3}^{\text{stoch}}$, was established in the proof of Theorem 1, as the structure of the martingale part is identical. Its limit is given by:
@@
The derivative of the stochastic part, as established in Corollary 1, yields the stochastic integral term. Summing these derivatives gives the full stochastic Volterra equation for $\dot{Y}_{t}$ as stated in (21).
@@
6 Future Work
Acknowledgement
Declarations
  • Author contribution: Yingli Wang led the research, authored the main manuscript, provided the majority of the proofs, prepared Figure 1, and developed Algorithm 1. Qinghua Wang contributed to the proof of Lemma 4 and played a key role in restructuring the manuscript. Ping He and Chunhao Cai supervised the project, provided critical revisions, and ensured the overall quality and coherence of the manuscript.
Funding Declarations

### v7

@@ Appendix B Proof of the Main Results
Finally, we turn to $T_{3}$. Remark that $M_{s}^{T}=\sqrt{\frac{T^{\alpha}\nu^{*}\delta^{-1}}{1-a_{T}}}Z_{s/T}^{T}$. Thus,
@@
Using Assumption 2.3, the prefactor converges to:
$\lim_{T\to\infty}\frac{1}{\sqrt{T^{\alpha}(1-a_{T})\nu^{*}\delta^{-1}}}=\frac{1}{\sqrt{\lambda\delta\cdot\nu^{*}\delta^{-1}}}=\frac{1}{\sqrt{\lambda\nu^{*}}}.$
Following the convergence arguments in Theorem 3.1 of Jaisson and Rosenbaum (2016), $T_{3}$ converges to
$\frac{1}{\sqrt{\nu^{*}\lambda}}\int_{0}^{t}\lambda(t-s)^{\alpha-1}E_{\alpha,\alpha}(-\lambda(t-s)^{\alpha})Z_{s}ds.$
Since $Z$ is a continuous martingale with quadratic variation $Y$, the fact that $Z_{t}=B_{Y_{t}}$ is a consequence of the Dambis-Dubin-Schwarz theorem (e.g., Theorem V-1.65 in Revuz and Yor (2013)).
B.3 Proof of Corollary 2.18
The proof extends the arguments used to establish Theorem 2.13. The core of the argument lies in analyzing the convergence of the scaled process $Y_{t}^{T}$ under the new time-dependent baseline intensity $\hat{\mu}_{T}(n)$.
@@
Let us denote the deterministic mean as $\bar{\Lambda}_{t}^{T}:=\mathbb{E}[Y_{t}^{T}]$. Using the expression for $\mathbb{E}[N_{n}^{T}]$ from Lemma 2.17 and applying the scaling factor, we can decompose $\bar{\Lambda}_{t}^{T}$ into several terms. After substituting the definition of $\hat{\mu}_{T}(n)$, the expression for $\mathbb{E}[N_{n}^{T}]$ becomes:
@@
The centered stochastic part, $Y_{t}^{T}-\bar{\Lambda}_{t}^{T}$, is obtained by scaling the expression from Equation (2.16) in Lemma 2.17. We denote this scaled stochastic part as $T_{3}^{\text{stoch}}$:
@@
  • By Assumption 2.3, it is straightforward to show that $\lim_{T\to\infty}T_{1}=0$.
  • For $T_{2}$, we use the same Riemann sum approximation argument as in the proof of Theorem 2.13, where the sum $\sum_{s=1}^{n-1}sA_{n-s}^{T}$ corresponds to the integral of the CDF. In the limit, the sum of the deterministic parts converges:
@@
  • The convergence of the scaled stochastic part, $T_{3}^{\text{stoch}}$, was established in the proof of Theorem 2.13, as the structure of the martingale part is identical. Its limit is given by:
@@
The derivative of the stochastic part, as established in Corollary 2.14, yields the stochastic integral term. Summing these derivatives gives the full stochastic Volterra equation for $\dot{Y}_{t}$ as stated in (2.17).